You rev your car's engine to 2800 rpm (rev/min).
Answer #1

a )

2800 revolutions per minute

1 revolutions in 1 / 2800 min

1 revolutions in 60 / 2800 sec = 0.0214 sec

then the time period of the engine is T = 0.0214 sec

b)

to calculate the frequency of the engine f is

f = 1 / time period of engine

then can be write as f = 1 / T

f = 1 / 0.0214

f = 46.72 Hz

c )

given value of time period is T = 0.043s

= 0.043 / 60 min

= 43 / 60000 min

= 1/43/60000

T = 1395.34 rev per min